# CMV-Fuse: Cross Modal-View Fusion of AMR, Syntax, and Knowledge Representations for Aspect Based Sentiment Analysis

### v1

@@ 6 Limitation
Appendix A Appendix A: Ablation study
A.1 Ablation Study
@@
A.1.1 Effects of Different Multi View Fusion Combinations
Table 3 systematically evaluates all combinations of our three core representations: AMR-based semantic ($H_{\text{amr}}$), syntactic structures ($H_{\text{syn}}$), and external knowledge ($H_{\text{kg}}$). Individual views show that $H_{\text{amr}}$ achieves the highest standalone performance (86.91% accuracy), validating semantic abstraction’s discriminative power for ABSA. Among pairwise combinations, $H_{\text{amr}}+H_{\text{kg}}$ (87.10%) performs best, demonstrating effective knowledge-semantic synergy, while $H_{\text{syn}}+H_{\text{kg}}$ (86.54%) shows minimal gains, indicating syntactic structures require semantic abstraction to effectively leverage external knowledge. The full three-way combination achieves optimal results (87.23%), confirming that systematic multi-view integration provides the most robust representation.
A.1.2 Effects of Different Multi View Loss Contributions
Table 4 reveals that contrastive learning is essential: removing $\mathcal{L}_{\text{amr}}$ causes the largest degradation (-0.94% accuracy, -2.88% F1), demonstrating critical importance of semantic consistency, while eliminating all contrastive objectives results in severe performance loss (-1.22% accuracy, -2.99% F1). Our full model with hierarchical fusion and all contrastive objectives achieves 87.76% accuracy and 81.99% F1, significantly outperforming the concatenation baseline (+0.53% accuracy).

### v2

@@ 6 Limitation
Appendix A Appendix A: Baseline Methods
We compare CMV-Fuse with several state-of-the-art models, including: (1) BERT Devlin et al. (2019), a general pre-trained model with a classification head for ABSA; (2) R-GAT Wang et al. (2020), which uses a relational GAT for aspect-based dependency structures; (3) DualGCN Li et al. (2021), a dual GCN that merges features from sentences and dependency trees; (4) dotGCN Chen et al. (2022), which uses a latent tree structure for aspects; and (5) SSEGCN Devlin et al. (2019), incorporates aspect-aware attention in GCNs.(6) KE-IGCN Wan et al. (2023) aims to select the highly relevant subgraphs, and proposes an interaction strategy to evaluate the interaction between external knowledge and the input text. (7) BiSyn-GAT Liang et al. (2022) graph-based model that captures syntactic dependencies in both bottom-up and top-down directions using attention mechanisms, enabling richer contextual representations for each token. (8) TextGT employs a double-view graph transformer to integrate syntactic and semantic dependencies for ABSA. (9) S2GSL models syntactic graphs with graph structure learning to improve aspect–opinion interaction. (10) Bert+CD enhances BERT with contextualized dependency(CD) representation, injecting syntactic dependency information into token embeddings. (11) DMAN, a Dual Mode Attention Network that jointly models semantic and syntactic information through complementary attention mechanisms. (12) MambaforGCN+Bert integrates the Mamba sequencing modeling architecture with GCN and BERT, leveraging long range dependency modeling from Mamba and structured syntactic reasoning from GCN.
Appendix B Appendix B: Ablation study
B.1 Ablation Study
We conduct comprehensive ablation studies to isolate and understand
  1. 1.
    the individual contribution of each view (AMR, DEP, CON, SEM, KG)
  2. 2.
    the effectiveness of complementary view pairings
  3. 3.
    the impact of different training objectives (e.g., alignment and contrastive losses) on performance gains.
  4. 4.
    The baselines considered included SOTA from mainly the last 3 years (Table 2)
This analysis clarifies which components drive performance gains and avoids attributing improvements to the framework without evidence (results in Appendix Table 3 and Table 4).
@@ B.1 Ablation Study
B.1.1 Effects of Different Multi View Fusion Combinations
Table 3 systematically evaluates all combinations of our three core representations: AMR-based semantic ($H_{\text{amr}}$), syntactic structures ($H_{\text{syn}}$), and external knowledge ($H_{\text{kg}}$). Individual views show that $H_{\text{amr}}$ achieves the highest standalone performance (86.91% accuracy), validating semantic abstraction’s discriminative power for ABSA. Among pairwise combinations, $H_{\text{amr}}+H_{\text{kg}}$ (87.10%) performs best, demonstrating effective knowledge-semantic synergy, while $H_{\text{syn}}+H_{\text{kg}}$ (86.54%) shows minimal gains, indicating syntactic structures require semantic abstraction to effectively leverage external knowledge.
For fair comparison, we concatenate different representations and feed them to a single-layer MLP classifier to achieve multi-view fusion. Note that we do not employ neither the proposed hierarchical fusion module nor any contrastive loss objectives to fuse the entire representation combination $H_{amr}$$H_{syn}$$H_{kg}$]; thus, the performances of the last row in Table 3 are slightly worse than the relative ones in Table 2. The full three-way combination achieves optimal results (87.23%), confirming that systematic multi-view integration provides the most robust representation.
@@ B.1.1 Effects of Different Multi View Fusion Combinations
B.1.2 Effects of Different Multi View Loss Contributions
Table 4 reveals that contrastive learning is essential: removing $\mathcal{L}_{\text{amr}}$ causes the largest degradation (-0.94 percentage points for accuracy and -2.88 percentage points for F1), demonstrating critical importance of semantic consistency, while eliminating all contrastive objectives results in severe performance loss (-0.38 percentage points for accuracy, -0.27 percentage points for F1). Our full model with hierarchical fusion and all contrastive objectives achieves 87.76% accuracy and 81.99% F1, significantly outperforming the concatenation baseline of 87.23% (+0.53 percentage points) accuracy and 80.82% (+1.17 percentage points) F1.